OpenStudy (anonymous):

A rectangular box is x feet long and x feet wide. The volume of the box is (4x^8+3x^6) cubic feet. What polynomial represents the height of the box in feet? A.4x+3 B.4x^6+3x^4 C.4x^8+3x^6 D.4x^7+3x^5

OpenStudy (anonymous):

V is L*W*h V= x*x*h =x^2*h = 4x^8 + 3x^6 so what must h be? (factor an x^2 from 4x^8 + 3x^6 )